# Should you use LLMs to simulate opinions? Quality checks for early-stage deliberation

### v1

@@ 3 Discussion
LLMs offer a compelling approach to simulating demographic opinions, particularly because they can generate plausible outputs without requiring any external data to get started. However, assessing whether these outputs faithfully capture the nuanced patterns of human perspectives across diverse groups, tasks, and topics remains a significant challenge. This difficulty creates space for premature or unsupported claims about their reliability. The problem is compounded by the fact that LLMs rarely refuse to answer; instead, they often produce responses that appear coherent and reasonable on the surface, even when the underlying logic or representational accuracy is flawed. In this paper, we propose a set of conceptual checks—focused on logical consistency, model stability, and alignment with basic stakeholder expectations—to provide a principled initial foundation for evaluating whether LLM-simulated opinions merit further investment and can inform prompt engineering efforts early in the research and development timeline.
Applying these methods to several state-of-the-art models, we provide empirical evidence that these models fail to yield reliable results across our three checks. These observed shortcomings can mislead decision-makers by distorting genuine opinion distributions and skewing outcomes in contexts such as market research, policy development, and strategic planning. Our findings underscore the need for caution and careful prompt engineering when deploying these technologies.
The proposed methodology serves as a practical framework for early-stage deliberation, equipping organizations with tools to assess LLM reliability before investing in large-scale human reference data. By integrating these quality checks, organizations can make more informed decisions about the deployment of LLMs and implement adequate risk management strategies. In the following sections, we analyze observed failure modes from our three experiments, their consequences, methodological implications, and provide insight into how organizations can use these tests to inform and manage the associated risk of these technologies.
@@
Experiment 1 reveals that most models fail to produce logically consistent neutral responses: although a neutral stance should logically be represented as a convex combination of the model’s predicted opinions for different demographic groups, only 20% of the 64 model–prompt–topic combinations passed this check. The remaining 80% produced “neutral” outputs that were actually more extreme than both men’s and women’s predicted opinions, casting doubt on whether these models reflect genuine opinion distributions or merely artifacts of their training. Such inconsistencies undermine the idea that opinion estimation might be a legitimate “emergent” LLM capability, thereby raising concerns about these models’ suitability for opinion surveys.
Methodological Implications
Logical consistency is a cornerstone of annotation quality: outputs that contradict each other without justification reduce trust in the annotator, whether human or machine. When LLMs violate basic structural expectations—such as the expectation highlighted in this experiment—it suggests a breakdown in internal reasoning. This failure challenges the use of LLMs as annotation tools or survey proxies. Our findings point to a critical gap: LLM evaluations need to account for internal relational consistency between related prompts, not just isolated outputs. Without this kind of internal logic check, high-quality-sounding responses may mask logical errors.
Risk Management
Although failing this check raises broader doubts about an LLM’s capacity to model group-level opinions, its practical impact depends on how central a coherent notion of neutrality is to the task at hand. If a use case primarily relies on modeling group-specific opinions, flawed neutral outputs may not be critical. However, for tasks like gauging overall sentiment or developing balanced recommendations, this failure mode becomes more pertinent. When these issues arise, multi-stage prompts that incorporate previously predicted group-level opinions as context might remedy them, allowing LLMs to position neutrality in their own context.
@@
Experiment 2 shows that even successive versions of the same model can exhibit unstable outputs, implying a lack of backwards compatibility of prompts. This instability complicates tasks such as policymaking, content moderation, and market research, where stable, interpretable opinions are critical. Third-party LLM APIs exacerbate the problem: providers frequently modify models to fix biases or improve performance, yet these opaque updates can invalidate prior workflows and expectations [50, 52].
Methodological Implications
Our findings highlight the importance of evaluating not just a single model, but also comparing different checkpoints (and, potentially, related models) to assess reliability and backwards compatibility. When outputs shift significantly between model versions, and there is no consistent sampling of human opinions over time, it becomes difficult to determine whether these changes reflect genuine shifts in public attitudes or are instead driven by exogenous factors, such as new training data, updated reinforcement learning objectives, or other changes in the model pipeline. This uncertainty poses a particular challenge for applications that rely on stability of decision-making criteria over time, such as content moderation. Moreover, unlike traditional software, LLMs often lack transparent, version-locked updates, reinforcing the need for continuous evaluation rather than one-time benchmarking.
Risk Management
Dedication to using a version-controlled model can help reduce some of the instability. Regardless, selecting the “best” model and prompt under these conditions is challenging: newer models are not guaranteed to outperform older ones since this task is certainly not directly optimized for, and improvements on one front (e.g., harmlessness or safety) may significantly impact prompt performance. While ensembling models can, in principle, mitigate uncertainty [53], shared training data and alignment constraints between models from the same model class may leave their biases too closely correlated to observe the usual benefits. Practitioners should test is whether in-context response calibration can offset the variability noted in our experiments. Future work should investigate which forms of contextual scaffolding—such as embedding representative edge cases or eliciting chain-of-thought reasoning—most effectively stabilize model outputs across checkpoints.
@@
Experiment 3 reveals systematic failures in how LLMs position group-level opinions. Older models like GPT-3.5-turbo often exaggerate opinion differences, even on common-sense, non-divisive claims, such as “An almond is a type of nut” and “A circle is round,” risking stereotyping and misallocation of resources in downstream applications like claim prioritization. Newer models (GPT-4, LLaMa3, Mistral, Titan), by contrast, frequently erode meaningful differences, particularly on topics known to be gender-divisive. This compression of diversity can obscure key priorities for certain groups. This may reflect an implicit refusal where models respond superficially but sidestep the task of representing nuanced disagreement.
We also observed inconsistent error direction across models. For the same topic (e.g., immigration), one model (e.g., GPT3.5a) may exaggerate differences while another (e.g. GPT-4) erodes them. These contradictions, combined with uncertainty in the human reference data itself, introduce irreducible uncertainty in model selection: we cannot definitively say which model is “closer to the truth” without assuming the reference data is unbiased or that one model is inherently more reliable than the other, both of which are challenging claims to validate in the context of opinion simulation.
Methodological Implications
Our approach—using curated topics in the TopicMisinfo dataset with strong societal priors—provides a domain-grounded alternative to noisy or expensive human reference data. If a model positions large gender gaps on statements like “A circle is round,” or erases known gender divides on issues like abortion, it likely fails to capture basic structures in group sentiment, and all of its outputs should be considered suspect. These tests serve a function analogous to inter-annotator agreement in human annotation pipelines [54, 26, 55]. Rather than treating human data as absolute ground truth, we assess whether models reflect well-established social patterns. This allows for lightweight, reproducible tests of model reliability without large-scale annotation efforts.
Risk Management
Organizations should prioritize testing whether model outputs align with clear expectations in their domain. Errors on divisive topics may reflect real-world complexity, but fabricating disagreement on neutral issues specifically designed for quality control can yield a high-quality signal that model performance is unreliable. In settings without established priors, domain experts can still construct targeted edge cases to test model reasoning. Including system instructions or information in the prompt that draws attention to the risks of both exaggeration and erosion of group differences may help calibrate responses with stakeholder expectations. Additionally, chain-of-thought prompts can be used to surface uncertainty or re-position group opinion estimation after an initial prediction [56].
@@ Conclusion
By applying quality control checks that surface logical consistency, LLM self-consistency over time, and stakeholder expectations of model performance, practitioners can determine whether an LLM clears a basic threshold for reliability before devoting significant resources to collecting large-scale human reference data to fine-tune model performance. A model that meets these conditions may still require subsequent testing and tuning, but it at least demonstrates fundamental viability. If it fails on multiple fronts, stakeholders should carefully consider whether an LLM-driven approach is worthwhile or whether alternative data-centric or human-led solutions would be more consistent and trustworthy.
@@ 4 Materials and Methods
All experiments in this study rely on a common bootstrap hypothesis-testing framework, following well-established guidelines [57, 58]. Although the test statistic and specific sampling procedures differ by experiment, the overall structure is as follows:
@@
    Define the test statistic. We specify a test statistic $L$ that measures the relevant difference or relationship of interest (e.g., the divergence between LLM and human responses, or between outputs from two LLM checkpoints).
@@
    Construct a pooled set under the null hypothesis. The null hypothesis $H_{0}$ varies by experiment but generally posits “no difference” in the phenomenon under investigation. Under $H_{0}$, we pool all annotations (or responses) across the conditions being compared, reflecting the assumption of no effect.
@@
    Bootstrap resampling. We generate $B=10{,}000$ bootstrap samples by resampling with replacement from the pooled set, preserving each condition’s original sample size. Each bootstrap sample yields a recalculated value of the test statistic, $L^{b}$.
@@
    Compute the p-value. Because all of our experiments use two-sided tests, the p-value is defined as the proportion of bootstrap replicates for which the absolute value of the bootstrapped statistic exceeds that of the observed statistic. Formally:
    $T=\frac{1}{B}\sum_{b=1}^{B}I\bigl{[}\bigl{|}\hat{L}^{b}\bigr{|}>\bigl{|}\hat{L}\bigr{|}\bigr{]},$ (1)
    where $\hat{L}$ is the observed (empirical) test statistic, and $\hat{L}^{b}$ is the statistic in the $b$-th bootstrap sample. We reject $H_{0}$ if $T<\alpha$, where $\alpha$ is the chosen significance level.
@@
$\hat{L}_{\omega}^{(b)}(q_{0})=\frac{1}{N_{\omega}}\sum_{i=1}^{N_{\omega}}\left|\hat{q}_{c_{i}}^{(b)}-q_{0}\right|,$
@@
indicating that the LLM’s neutral responses for topic $\omega$ can be exactly recovered by a convex combination with weight $q_{0}$. We reject $H_{0}$ if $\hat{L}_{\omega}(q_{0})$ is unusually large relative to its bootstrap distribution. We repeat this procedure for $q_{0}$ values in increments of 0.05 over $[0,1]$.
@@
$\hat{L}_{\omega}^{j-k}=\frac{1}{N_{\omega}}\sum_{i=1}^{N_{\omega}}\hat{L}_{c_{i}}^{j-k}.$
@@
$\hat{L}_{\omega}=\frac{1}{N_{\omega}}\sum_{i=1}^{N_{\omega}}\hat{L}_{c_{i}}.$
@@
$\hat{L}_{\omega}=\frac{1}{N_{\omega}}\sum_{i=1}^{N_{\omega}}\hat{L}_{c_{i}}.$
@@ Acknowledgments
This chart complements the results from Experiment 3, as seen in Table 2, by showing the positioning of LLM-generated opinions on a Likert scale for men and women.
@@
S5: Experiment 1 Replication with Conditional Prompt 2
As a robustness check, we replicated Experiment 1 except we changed the prompt used to gather LLM annotations from Conditional Prompt 1 to Conditional Prompt 2. The main difference between these prompts is the Prompt 1 mentions gender twice while Prompt 2 mentions gender only once. While the individual values of $\hat{E}_{\omega}$ may be different than Experiment 1, the overall results appear qualitatively very similar to those obtained from the experiment. The gpt-3.5-turbo class of models tend to exaggerate differences in less divisive topics, while the Mistral-Large and Titan-Text-Premier significantly eroded differences that actually exist among genders.
@@
Abortion 0.25 ** -0.54 *** -0.26 * -0.65 *** -0.50 ***
Black Americans 0.26 * -0.42 ** -0.10 -0.60 *** -0.38 **
Illegal Immigration 0.20 -0.47 *** 0.19 -0.46 *** -0.49 ***
LGBTQ 0.43 *** -0.28 * -0.29 * -0.38 *** -0.43 ***
@@
Entertainment 0.26 0.32 * 0.56 *** 0.26 ** 0.12
Gold 0.09 0.43 *** 0.17 ** -0.00 -0.03
HealthScience 0.41 *** -0.04 0.66 *** -0.02 -0.08
Sports 0.02 0.11 0.03 -0.16 -0.10
USA 0.71 *** 0.09 0.40 *** -0.06 -0.16 **
WeatherClimate 0.09 0.28 0.41 * 0.12 -0.12
@@
Table 10: Results: Replication of Experiment 1 with Conditional Prompt 2

### v2

@@ 3 Discussion
LLMs offer a compelling approach to simulating demographic opinions, particularly because they can generate seemingly plausible outputs without requiring any additional data to get started. However, assessing whether these outputs faithfully capture the nuanced patterns of human perspectives across diverse groups, tasks, and topics remains a significant challenge. This difficulty creates space for premature or unsupported claims about their reliability. The problem is compounded by the fact that LLMs rarely refuse to answer; instead, they often produce responses that appear coherent and reasonable on the surface, even when the underlying logic or representational accuracy is flawed. In this paper, we propose a set of conceptual quality checks—focused on logical consistency, prompt stability, and alignment with basic stakeholder expectations—to provide a principled initial foundation for evaluating whether LLM-simulated opinions merit further investment. The proposed tests can inform prompt engineering efforts early in the research and development timeline.
Applying these quality checks to several state-of-the-art models, we provide empirical evidence that prompting these models with a common and intuitive approach to simulating opinions—“backstory” prompting [15]—fails to yield reliable results across our three checks. These observed shortcomings can mislead decision-makers in contexts such as market research, policy development, and strategic planning. Our findings underscore the need for caution and careful prompt engineering when deploying these technologies.
The proposed methodology serves as a practical framework for early-stage deliberation, equipping organizations with tools to assess LLM reliability before investing in large-scale human reference data. By integrating these quality checks, organizations can make more informed decisions about the deployment of LLMs and implement adequate risk management strategies. In the following sections, we analyze observed LLM failure modes revealed by our three tests and their consequences, followed by an assessment of how organizations can use these tests to inform and manage the associated risk of these technologies for opinion simulation.
@@
Experiment 1 reveals that most models fail to produce logically consistent neutral responses: although a neutral stance should logically (and statistically) be represented as a convex combination of the model’s predicted opinions for different demographic groups, only 20% of the 64 model–prompt–topic combinations passed this check. The remaining 80% produced “neutral” outputs that were actually more extreme than both men’s and women’s predicted opinions.
Notice the difference between the type of logical inconsistency at issue here with other cases of inconsistency previously observed in the literature. Most earlier studies focus on sentence-level or reasoning-chain contradictions—for example, mutually incompatible factual beliefs [51], self-inconsistent chain-of-thought traces [52], violations of basic propositional logic [53], or within-passage contradictions [54]. By contrast, the inconsistency we expose appears at the distributional level: the joint set of simulated opinions cannot all be true because the neutral distribution lies outside the feasible region defined by the demographic-specific distributions.
While one might dismiss sentence- or reasoning-level logical errors by noting that people themselves reason inconsistently, that defense fails here. If averaging men’s and women’s simulated opinions cannot approximately reproduce the model’s “neutral” opinion, then at least one of the three distributions is necessarily wrong. Because these distributions are meant to substitute for human survey data, their internal incompatibility undermines any downstream use. Without further insight into which distribution is incorrect, a practitioner cannot simply acknowledge the issue and proceed; they must either alter their simulation approach or abandon the simulated data altogether.
Risk Management Implications
Logical consistency is a cornerstone of subjective annotation quality: outputs that contradict each other without justification reduce perceived annotator reliability, whether human or machine [32]. The stark misalignment between “neutral” and subgroup opinions suggests that LLM assessment of simulated opinions must move beyond scoring isolated answers and instead evaluate the internal algebra that should bind related outputs from different demographic distributions. This implies the need for new relational benchmarks that probe whether LLMs maintain coherent beliefs across prompted demographic conditions. Such benchmarks would explicitly test algebraic relationships—e.g., requiring the “neutral” distribution to fall within the convex hull of subgroup predictions—and can be generalized to any setting where aggregation or probabilistic constraints link multiple model outputs. While our study examines differences between men and women, relational benchmarks should incorporate broader relationships from numerous demographic categories relevant to organizations, including intersectional characteristics. In such cases, it would be necessary to show that neutrality holds along each demographic axis and their intersection.
Existing evidence shows that a related class of logical inconsistencies—intransitive preferences—persist even after adding chain-of-thought prompts or external knowledge, indicating that post-hoc “knowledge injection" alone cannot guarantee coherence [55]. A more principled remedy may be to embed the desired structure inside the learning objective itself. Recent research has operationalized this idea through “semantic loss", a differentiable penalty that forces a network’s outputs to satisfy symbolic rules, including total-order rankings and simple-path constraints [56]. Similarly, one could encode algebraic relationships among simulated opinion distributions and optimize the model under those constraints. For this, it should be noted that directly optimizing for the proposed quality test would be undesirable and could yield an instantiation of Goodhart’s Law [57], by which making a measure the target erodes its value. Instead, it would be necessary to develop a complete set of algebraic constraints that are not only necessary but also sufficient to ensure the quality of simulated opinions.
@@
Experiment 2 shows that even successive versions of the same model can exhibit unstable opinion outputs holding the prompt constant, implying a lack of backwards compatibility of prompt outputs. This instability complicates tasks such as policymaking, content moderation, and market research, where, all else equal, stable, interpretable opinions are critical. Third-party LLM APIs exacerbate the problem: providers frequently modify models to fix biases or improve performance, yet these opaque updates can invalidate prior workflows and expectations [49, 58].
Risk Management Implications
Our findings highlight the importance of evaluating models not just as isolated, static entities, but also comparing different checkpoints (and, potentially, related models) to assess reliability and backwards compatibility of outputs. When outputs shift significantly between model versions, and there is no consistent sampling of human opinions over time, it becomes difficult to determine whether these changes reflect genuine shifts in public attitudes or are instead driven by exogenous factors, such as a shift in the distribution of the training data, a change in reinforcement learning objectives, modified safeguards, or other changes in the model pipeline. This uncertainty poses a particular challenge for applications that rely on stability of decision-making criteria over time, such as content moderation. Moreover, unlike traditional software, LLMs often lack transparent, version-locked updates, reinforcing the need for continuous evaluation rather than one-time benchmarking.
Dedication to using a version-controlled model can help reduce some of the instability. Regardless, selecting the “best” model and prompt under these conditions is challenging: newer models are not guaranteed to outperform older ones since opinion simulation is certainly not directly optimized for, and improvements on one front (e.g., harmlessness or safety) may significantly impact prompt performance [49]. For example, a guardrail meant to mitigate concerns of representational bias could inadvertently erode a model’s ability to simulate gender differences in opinion. While ensembling models can, in principle, mitigate uncertainty [59], shared training data and alignment constraints between models from the same model class may leave their biases too closely correlated to observe the usual benefits.
We evaluated our proposed quality checks with the widely studied backstory prompting technique because of its prominence in recent work. However, the same diagnostics can just as readily assess whether alternative strategies—such as in-context response calibration—successfully mitigate the variability we observed. By applying the checks across methods, researchers gain a consistent yardstick for comparing not only models but also the prompting paradigms that steer them.
@@
Experiment 3 reveals systematic failures in how LLMs position group-level opinions. Older models like GPT-3.5-turbo often exaggerate opinion differences, even on common-sense, non-divisive claims, such as “An almond is a type of nut” and “A circle is round,” risking stereotyping, misrepresenting and misallocation of resources in downstream applications like claim prioritization. Newer models (GPT-4, LLaMa3, Mistral, Titan), by contrast, frequently erode meaningful differences, particularly on topics known to be gender-divisive. This erosion of diversity can obscure key priorities for certain groups. This may reflect an implicit refusal where models respond to the prompt but refuse to display disagreement across genders.
Risk Management Implications
Organizations should prioritize testing whether model outputs align with clear expectations in their domain, and they should pay close attention to errors that contradict their priors. For instance, an error of exaggeration of differences between groups on a non-divisive issue is likely a more serious error than an error of exaggeration on an issue known to be divisive. In settings where established priors on special topics are scarce, domain experts can still construct targeted edge cases to test model reliability.
Notice that the types of errors we saw in Experiment 3a were correlated to the errors we saw in Experiment 3b. That is, models that predicted statistical differences between genders in “Gold” claims in Experiment 3a tended to also exaggerate gender differences relative to human data, especially on non-divisive topics, in Experiment 3b. Likewise, models that showed no difference between genders for claims in the topic “Abortion” in Experiment 3a tended to erode differences on divisive topics relative to humans in Experiment 3b. This shows that the simple approach of testing edge cases with known priors without human data can yield a plausible signal for assessing simulated opinion alignment with stakeholder expectations.
Moreover, while our current metrics focus on mean differences in opinion distributions, some stakeholders may prioritize other distributional properties—such as variance or skew [18, 17]. Our bootstrap hypothesis testing framework is flexible: the test statistics we propose can be adapted to target other distributional moments, supporting a wider range of stakeholder needs and use cases.
@@ Conclusion
By applying quality control checks that surface logical consistency, prompt stability over time, and alignment with stakeholder domain knowledge, practitioners can determine whether an approach to opinion simulation clears a basic threshold for reliability before devoting significant resources to collecting large-scale human reference data. An approach that meets these conditions may still require subsequent testing and tuning, but it at least demonstrates fundamental viability. If it fails on multiple fronts, stakeholders should carefully consider whether an LLM-driven approach is worthwhile or whether alternative data-centric or human-led solutions would be more consistent and trustworthy.
@@ 4 Materials and Methods
All experiments rely on a shared bootstrap hypothesis-testing framework following established guidelines [60, 61]. Although specific test statistics vary by experiment, the overall procedure is:
@@
    Define the test statistic. We specify a statistic $L$ capturing the quality check.
@@
    Construct pooled set under $H_{0}$. This pooled set, which we refer as $\mathcal{D}$, reflects the assumed distribution from which the observed test statistic would be drawn if the quality check is satisfied.
@@
    Bootstrap resampling. Generate $B=10{,}000$ bootstrap samples by resampling with replacement from $\mathcal{D}$, matching the original sample size observed in the data. Each bootstrap sample provides a recalculated statistic $L^{b}$.
@@
    Compute p-value. For two-sided tests, the p-value is the proportion of bootstrap samples whose statistic’s absolute value exceeds the observed statistic:
    $T=\frac{1}{B}\sum_{b=1}^{B}I\bigl{[}\bigl{|}\hat{L}^{b}\bigr{|}>\bigl{|}\hat{L}\bigr{|}\bigr{]}.$ (1)
    Reject $H_{0}$ if $T<\alpha$, where $\alpha$ is the significance threshold.
@@
$\hat{L}_{\omega}^{(b)}(q_{0})=\frac{1}{N_{\omega}}\sum_{i=1}^{N_{\omega}}\left|\hat{q}_{c_{i}}^{(b)}-q_{0}\right|,$ (7)
@@
indicating that the LLM’s neutral responses for topic $\omega$ can be exactly recovered by a convex combination with weight $q_{0}$. We reject $H_{0}$ if $\hat{L}_{\omega}(q_{0})$ is unusually large relative to its bootstrap distribution, as in Equation 1. We repeat this procedure for $q_{0}$ values in increments of 0.05 over $[0,1]$. We correct for multiple hypothesis testing by setting $\alpha^{*}=0.05/20=0.0025$.
@@
$\hat{L}_{\omega}^{j-k}=\frac{1}{N_{\omega}}\sum_{i=1}^{N_{\omega}}\hat{L}_{c_{i}}^{j-k}$ (9)
@@
$\hat{L}_{\omega}=\frac{1}{N_{\omega}}\sum_{i=1}^{N_{\omega}}\hat{L}_{c_{i}}.$ (11)
@@
$\hat{L}_{\omega}=\frac{1}{N_{\omega}}\sum_{i=1}^{N_{\omega}}\hat{L}_{c_{i}}.$ (13)
@@ Acknowledgments
S0: Literature Review Table
Below is a table, grouped by outcome, that summarizes the literature on whether LLMs are effective at simulating group-level opinions. These studies have generally used small variations of "backstory" prompting, but there are also some that examine fine-tuning and in-context learning approaches, as well as other forms of prompt engineering.
Table 3: Overview of studies using LLMs to simulate human opinions, grouped by outcome
Title Task Approach Dataset LLMs Key Findings
Out of One, Many [15] Political Attitudes Prompt Engineering: Backstories ANES GPT-3 Positive. Useful for simulating opinion distributions.
Donald Trumps [19] Voting Prompt Engineering: Backstories + structural ANES, WVS ChatGPT Positive. Backstory prompts produce outputs that match voting patterns.
Synthesizing Public Opinion [21] Political Attitudes In-Context Learning: Knowledge Injection CES Gemma2, LLaMa3 Positive. Incorporating knowledge in prompts improves over backstory.
Mixture of Personas [18] Sentiment Prompt Engineering: Probabilistic Mixture of Personas IMDB, Yelp, AGNews LLaMa3 Positive. Better alignment and within-group diversity than backstory prompting.
Modeling Targeted Populations [20] EV attitudes Fine-tuning: ID & Demographics BEV Survey LLaMa2 Positive. Fine-tuning improved KL-divergence and RMSE with subgroups.
Aligning with Whom? [22] Politeness Prompt Engineering: Backstories POPQUORN GPT-3.5, GPT-4 Mixed. Aligns more with White participants than Black/Asian ones.
Whose Opinions? [9] Political Attitudes Prompt Engineering: Backstories OpinionQA GPT-3.5 Mixed. Matches younger/liberal demos; some steering helps.
Climate Opinion [26] Environmental attitudes Prompt Engineering: Backstories + covariates Novel. Nat’l sample GPT-3.5, GPT-4 Mixed. Demographics alone insufficient; psych vars improve results.
Synthetic Survey Data [25] Feeling thermometers Prompt Engineering: Backstories ANES GPT-3.5-Turbo Mixed. Similar means, reduced variance; prompt/time sensitive.
Beyond Demographics [24] Content moderation Fine-tuning: IDs & Demographics MINT, POPQUORN, DICES LLaMa 3 Mixed. Models learn individual opinions, but fail at demographic-level.
Chameleons? [23] Political Attitudes Prompt Engineering: Backstories ESS GPT-3.5, Llama 2/3, Mistral, Deepseek Negative. Highly sensitive to prompt; misrepresents group-level opinions.
LLMs Flatten Identities [17] Political Attitudes Prompt Engineering: Backstories Novel (n=3200) GPT-4, Llama 2, WizardLM Negative. Misrepresents minorities; flattens within-group variation.
Cross-Cultural Simulation [6] Political Attitudes Prompt Engineering: Backstories WVS GPT-3.5-Turbo Negative. Performs better on U.S. samples; shows liberal bias.
@@
This chart complements the results from Experiment 3, as seen in Tables 1 and 2, by showing the positioning of LLM-generated opinions on a Likert scale for men and women.
@@
S5: Salient Examples of Opinion Differences Diverging from Priors in Exp 3a
Below, we highlight several claim-level differences worth noting that deviate from our strong priors on how opinions should be ordered.
Table 11: $\mu(c_{i})^{gender}_{LLM}$ represents the average predicted harmfulness rating for the claim for a given gender.
S6: Experiment 3b Replication with Conditional Prompt 2
As a robustness check, we replicated Experiment 1 except we changed the prompt used to gather LLM annotations from Conditional Prompt 1 to Conditional Prompt 2. The main difference between these prompts is the Prompt 1 mentions gender twice while Prompt 2 mentions gender only once. While the individual values of $\hat{L}_{\omega}$ may be different than Experiment 1, the overall results appear qualitatively very similar to those obtained from the experiment. The gpt-3.5-turbo class of models tend to exaggerate differences in less divisive topics, while the Mistral-Large and Titan-Text-Premier significantly eroded differences that actually exist among genders.
Table 12: Combined Results for Multiple Models
@@
Abortion 0.42 ** -0.77 *** -0.41 * -0.90 *** -0.70 ***
Black Americans 0.47 ** -0.55 * -0.08 -0.76 *** -0.47 **
Illegal Immigration 0.28 -0.68 *** 0.33 -0.65 *** -0.70 ***
LGBTQ 0.64 *** -0.35 -0.45 * -0.53 ** -0.61 ***
@@
Entertainment 0.32 0.41 * 0.75 *** 0.30 ** 0.14
Gold 0.12 0.60 *** 0.24 ** -0.00 -0.03
HealthScience 0.49 *** -0.06 0.88 *** -0.03 -0.09
Sports 0.02 0.16 0.09 -0.22 -0.13
USA 0.96 *** 0.17 0.55 *** -0.08 -0.20 *
WeatherClimate 0.15 0.39 0.56 * 0.16 -0.17